how do i place fractions on a number line when the fractions arnt the same, Example: -4 5/6, -4 3/7 ??
Answer by checkley77(12844)

Use the decimal values.
-4 5/6=-(4*6+5)/6=-(24+5)/6=-29/6=-4.8333
-4 3/7=*-(4*7+3)/7=-(28+3)/7=-31/7=-4.4286